# Lessons Learned from a Unifying Empirical Study of Parameter-Efficient Transfer Learning (PETL) in Visual Recognition

### v2

@@ 3 How Do PETL Methods Perform in Low-Shots Regime?
Table 1: Results on VTAB-1K (19 tasks from 3 groups). Based on the accuracy among PETL, linear probing, and full fine-tuning, we find two task groups (purple and orange), as discussed in section 6.
@@
Figure 2: Ranking frequency of 15 methods (14 PETL + linear probing) for three groups in VTAB-1K. Element $(i,j)$ is the number of times method $i$ ranks $j^{th}$ in each group. Methods are ordered by mean ranks (in brackets). The parameters column shows the # of trainable parameters in millions. More details are in Appendix C.
@@
Pre-trained models are meant to ease downstream applications. One representative scenario is low-shot learning: supervised fine-tuning with a small number of examples per class. Indeed, low-shot learning has been widely used to evaluate PETL performance.
@@
Dataset. VTAB-1K (Zhai et al., 2019) consists of 19 image classification tasks from three groups. The Natural group comprises natural images captured with standard cameras. The Specialized group contains images captured by specialist equipment for remote sensing and medical purposes. The Structured group evaluates the scene structure comprehension, such as object counting and 3D depth estimation. Following Zhai et al. (2019), we perform an 80/20 split on the 1000 training images in each task for hyperparameter tuning. The reported top1 accuracy is obtained by training on the 1000 training images and evaluating on the original test set.
@@
Methods. We consider linear probing, full fine-tuning, and 14 PETL methods including 2 prompt-based (Jia et al., 2022), 5 adapter-based (Houlsby et al., 2019; Pfeiffer et al., 2021; Chen et al., 2022b; Jie & Deng, 2022; Luo et al., 2023), 4 Direct selective (Zaken et al., 2022; Basu et al., 2023; Xie et al., 2023; Lian et al., 2022), and 3 Efficient selective (Hu et al., 2021; Jie & Deng, 2023). Please refer to subsection 2.2 for details.
@@
Setup. We employ the ViT-B/16 model (Dosovitskiy et al., 2020) pre-trained on ImageNet-21K (Deng et al., 2009) as the backbone. The prediction head is randomly initialized for each dataset. Images are resized to $224\times 224$. We systematically tune 1) learning rate, 2) weight decay, and 3) approach-specifics like the size of PETL parameters which are often left intact. We set a cap for 3), $\leq 1.5\%$ of ViT-B/16. We also turn the drop path rate (Huang et al., 2016) on (e.g., 0.1) or off (i.e., 0). A detailed hyperparameter search grid and additional training details are provided in Appendix A.1.
@@
Results. As shown in 1(a) and Table 1, PETL methods generally outperform both linear probing and full fine-tuning across datasets. Additionally, under fair hyper-parameter tuning, we surprisingly found that most PETL methods perform similarly as the relative standard deviations (divided by the means) in all three groups are quite low. Simple methods (e.g., Bitfit) and PETL methods originally proposed for NLP (e.g., LoRA and Adapter), which were previously reported as inferior due to un-optimized implementations and hyperparameter tuning, now demonstrate competitive performance with SOTA visual PETL methods. To understand the relative advantages of different approaches, we provide the ranking frequency of PETL methods across different groups in Figure 2, where the element $(i,j)$ in each ranking matrix represents the frequency that method $i$ ranks $j^{th}$ in each group. Methods are ordered by their mean ranks (in brackets), and the parameters column indicates the number of trainable parameters in millions. In the natural group, simpler methods with fewer trainable parameters—such as DiffFit and Fact-TT—offer a cost-effective solution without compromising performance. Conversely, in the specialized and structured groups, methods with more parameters generally yield better performance. We hypothesize that this performance discrepancy arises from the domain affinity between the pre-trained domain (ImageNet) and the downstream domains. The natural group, sharing a stronger affinity with ImageNet, allows simpler methods like BitFit to adjust the features effectively. In contrast, the specialized and structured groups necessitate more complex methods with more trainable parameters to bridge the domain gap.
@@
Recipes. In low-shot regimes, when the downstream data are similar to the pre-trained data, simple methods (e.g., DiffFit) with decent accuracy and much fewer parameters are preferred—aligning with Occam’s razor. When there is a substantial domain gap, complex methods with higher accuracy (but more parameters) become competitive. Learning with low-shot data is prone to over-fitting. We find that if the drop path rate — which stochastically drops a transformer block per sample (Huang et al., 2016) — is set not as default (i.e., nonzero), all the methods can benefit from such a regularization, as shown in Figure 10 in the Appendix.
4 PETL Approaches Offer Complementary Information
@@
Figure 3: Prediction similarity analysis: element $(i,j)$ shows the percentage of samples that method $i$ and method $j$ have the same predictions. Although different methods achieve similar accuracy, they have diverse predictions. Details in Appendix C.
Figure 4: LoRA, Adapter and SSF fine-tuned on CIFAR100(VTAB-1K). (a): correct prediction overlap for the 5K most confident data. (b): wrong prediction overlap for the 5K least confident data. Details in Appendix C.
The previous section demonstrates that all PETL methods perform similarly across various domains. Given that different PETL methods are trained on the same downstream data using the same backbone and achieve comparable accuracy, one might expect them to learn similar knowledge from the data, resulting in similar predictions. Contrary to this expectation, our findings below reveal that different PETL methods acquire distinct and complementary knowledge from the same downstream data, even when built upon the same backbone, leading to diverse predictions.
Figure 5: Ensemble (avg logits) provides consistent gain on most datasets thanks to the diverse prediction. Details in Appendix C.
We start by analyzing their prediction similarity on the same dataset in VTAB-1K. It is expected that their predictions are similar for datasets with very high accuracy, such as Flowers102 (avg 99.1%) and Caltech101 (avg 91.4%). Beyond them, we find that most PETL methods show diverse predictions in other datasets in VTAB-1K. Figure 4 shows the prediction similarities between 14 PETL methods in DTD, Retinopathy, and DMLab, which belong to natural, specialized, and structured groups, respectively. In DTD and Retinopathy, most methods differ in about 20% of their predictions, while in DMLab, this difference increases to approximately 35%, even though they achieve similar accuracies. This prediction diversity may be attributed to the different inductive biases (Neyshabur et al., 2014) of PETL methods — they explicitly select specific parameters to update or insert different modules at various locations within the model. More analyses and details about Figure 4 are offered in Appendix C.
Such diverse predictions across methods open up the possibility of leveraging their heterogeneity for further improvement. The most straightforward approach is ensemble (Gontijo-Lopes et al., 2021), e.g., average logits over methods. Figure 5 demonstrates the ensemble performance gain over all the PETL methods in each dataset, where we use the worst PETL method as the baseline. Due to the diverse predictions across methods, the ensemble can provide consistent gain.
Also, we analyze if PETL methods make similar correct predictions for high-confidence samples and similar mistakes for low-confidence samples. Figure 4 shows the correct prediction overlap for the 5K most confident samples (per method) and the wrong prediction overlap for the 5K least confident samples (per method). For demonstration purposes, we select one method from each PETL category (LoRA, Adapter, SSF) and they are fine-tuned on CIFAR-100 in VTAB-1K. Methods within the same category also show diverse predictions (Appendix C). Since they make different predictions in both high and low-confidence regimes, this paves the way for new possibilities of using different PETL methods to generate diverse pseudo-labels for semi-supervised learning Yang et al. (2022); Gan & Wei ; Wang et al. (2022b), domain adaptation Farahani et al. (2021); Tang et al. (2024); Englert et al. (2024), continual learning Shim et al. (2021); Lomonaco et al. (2022); Mai et al. (2021) and transfre learning Zhuang et al. (2020); Tu et al. (2024); Liu et al. (2022c).
5 How Do PETL Methods Perform in Many-Shot Regime?
Recent works in NLP Chen et al. (2022a) have indicated that PETL methods may not perform as competitively as full fine-tuning when data is abundant. We thus aim to investigate PETL’s performance in many-shot regimes by addressing the following questions: (1) Should we use PETL or full fine-tuning when data is sufficient? (2) How should we adjust the number of trainable parameters for PETL methods in many-shot regimes?
Dataset. We select one representative dataset from each of the natural, specialized, and structured groups in VTAB: (1) CIFAR-100 Krizhevsky et al. (2009), a natural image dataset comprising 50K training images across 100 classes; (2) RESISC Cheng et al. (2017), a remote sensing dataset for scene classification with 25.2K training samples across 45 classes; and (3) Clevr-Distance Zhai et al. (2019); Johnson et al. (2017), a synthetic image dataset for predicting the depth of the closest object from the camera with 6 depth classes and 70K samples. The reported results are obtained by training on the full training set and evaluating on the original test set.
Setup. The model setup follows the VTAB-1K experiment. More details about setup and hyperparameter search are provided in Appendix A.
Results. In many-shot regimes, with sufficient downstream data, full fine-tuning may catch up and eventually outperform PETL methods. However, from Figure 6, we found that even in many-shot regimes, PETL can achieve comparable results with full fine-tuning, even just using $2\%$ of fine-tuning parameters. (The performance gain, however, quickly diminishes and plateaus after 5% of tunable parameters.) By comparing the results on the domain-close CIFAR-100 and domain-different RESISC and Clevr, we have some further observations. On the one hand, downstream tasks with larger domain gaps suggest the need to update, perhaps many, parameters to obtain high accuracy. With sufficient downstream data, full fine-tuning is less prone to over-fitting and indeed attains a high accuracy. But interestingly, PETL methods, with only $2\sim 5\%$ of tunable parameters, achieve similar accuracy, suggesting that its design principle does offer sufficient effective capacity for the model to learn Zhang et al. (2021). On the other hand, downstream tasks with smaller domain gaps suggest that the pre-trained model had learned sufficient knowledge about them; fully fine-tuning it thus risks washing such knowledge away. In fact, we found that PETL notably outperforms full fine-tuning on CIFAR-100, suggesting it as a more robust transfer learning algorithm for downstream tasks.
Recipes. In many-shot regimes, PETL methods with sufficient parameters ($2\sim 5\%$) appear more favorable than full fine-tuning and linear probing. On the one hand, they achieve comparable and even better accuracy than full fine-tuning. On the other hand, the tunable parameters remain manageable. The parameter efficiency of PETL also often implies less training GPU memory usage and training time, making PETL methods a favorable alternative in many-shot regimes. For a downstream domain that is close to the pre-training domain, PETL shows much pronounced transferability. For a downstream domain that is quite different, the limited tunable parameters (controversially, $2\sim 5\%$ already amounts to a few million) already allow the model to learn sufficiently.
Refer to caption Refer to caption Refer to caption
@@ 5 How Do PETL Methods Perform in Many-Shot Regime?
Figure 6: PETL accuracy in many-shot regimes, with different parameter sizes (X-axis) on three datasets from different domains. Even 2%-5% trainable parameters allow the models to have sufficient capacity to learn from full data. (Details are in Appendix C)
6 Why Do PETL Methods Work?  [^2] Our intention is not to offer a definitive conclusion about why PETL works. As discussed in subsection 2.3, there is currently no universally agreed-upon explanation for the effectiveness of PETL. We hope our empirical findings will contribute to the ongoing efforts to understand the underlying principles of PETL methods.
@@
Putting together section 3 and section 5, we identify two distinct patterns regarding the performance among linear probing, full fine-tuning, and PETL. Within 19 VTAB-1K tasks, we see: (1) Full fine-tuning outperforms linear probing. As linear probing reflects the pre-trained feature quality for downstream tasks, case (1) suggests the necessity to update the backbone to close the gap between pre-trained and downstream domains. (2) Linear probing surpasses full fine-tuning, suggesting the pre-trained features are good enough (at least in a low-shot scenario). Recklessly updating them may risk over-fitting. Figure 7 (a-b) summarizes the low-shot accuracy comparison based on the categorization above; each line corresponds to one task. Linear probing, PETL, and fine-tuning are located in order, from left to right, to reflect their tunable parameter sizes. PETL’s superiority in both cases showcases its capacity to learn and its regularization role to prevent over-fitting.
@@
We also draw the many-shot accuracy in Figure 7 (c-d) based on the same categorization: RESISC and Clevr in case (1), and CIFAR-100 in case (2). In the many-shot setting, full fine-tuning consistently outperforms linear probing, which seems to suggest no more risk of over-fitting. However, on CIFAR-100 (Figure 7 (d)), we again see a noticeable gap between PETL and full fine-tuning, just like in Figure 7 (b). Such a concave shape reminds us of the long-standing under-fitting-over-fitting curve, suggesting that even with sufficient downstream data, full fine-tuning still risks over-fitting.
@@
Taking into account PETL’s comparable performance to full fine-tuning on RESISC and Clevr with large domain gaps, we conclude — PETL succeeds as a high-capacity learner equipped with an effective regularizer; the two roles trade-offs well such that PETL can excel in both low-affinity and high-affinity domains under both low-shot and many-shot settings.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d)
@@
Figure 7: (a): VTAB-1K tasks in case (1) , PETL $>$ full $>$ linear . (b) VTAB-1K tasks in case (2), PETL $>$ linear $>$ full. (c) RESISC & Clevr in case (1) with enough data, PETL $\approx$ full $>$ linear. (d) CIFAR in case (2) with enough data, PETL $>$ full $>$ linear. Within each figure, left for linear, middle for PETL, and right for full. More details are in Appendix C.
7 Are PETL methods more robust to distribution shifts?
Large pre-trained models such as CLIP Radford et al. (2021) and ALIGN Jia et al. (2021) have demonstrated unprecedented accuracy across a range of data distributions when performing zero-shot inference. However, recent studies Wortsman et al. (2022); Radford et al. (2021) have shown that fine-tuning on downstream data, while significantly boosting performance on the target distribution, often compromises the model’s robustness to distribution shifts. Given that PETL only updates a limited number of parameters in the model, we investigate whether PETL can offer a more robust alternative to full fine-tuning for pre-trained models.
Dataset. We use 100-shot ImageNet-1K as our target distribution, with each class containing 100 images. Following Wortsman et al. (2022), we consider 4 natural distribution shifts from ImageNet: ImageNet-V2 Recht et al. (2019), a new ImageNet test set collected with the original labeling protocol; ImageNet-R Hendrycks et al. (2021a), renditions for 200 ImageNet classes; ImageNet-S Gao et al. (2023), sketch images for 1K ImageNet classes; ImageNet-A Hendrycks et al. (2021b), a test set of natural images misclassified by a ResNet-50 He et al. (2015) for 200 ImageNet classes.
Setup. We focus on the CLIP ViT-B/16 model, which comprises a visual encoder and a text encoder, pre-trained via contrastive learning on image-text pairs. Following Wortsman et al. (2022), we add an FC layer as the prediction head with zero-initialized bias and initialize weights using the class label text embedded by the text encoder. Subsequently, we discard the text encoder and apply PETL methods to the visual encoder, fine-tuning only the PETL modules and the head. More details about the CLIP model and experiment setup can be found in Appendix A.1.
Results. As shown in Table 2, while some PETL methods may not surpass full fine-tuning on the target distribution, they consistently demonstrate more robust performance on distribution shift data. This is likely because PETL updates only a small fraction of the parameters, thus preserving the robust features of the foundation models. Given the similar target distribution performance, should we blindly use PETL methods for more robustness?
Weight-space ensembles (WiSE) for PETL. WiSE (Wortsman et al., 2022), which linearly interpolates the full fine-tuned and original models, is a popular fine-tuning approach to enhance robustness. We explore whether WiSE can enhance the robustness of PETL. To apply WiSE to PETL, we first linearly interpolate the prediction head with a mixing coefficient $\alpha$. For direct selective tuning methods (e.g.BitFit), we directly interpolate with the original model. Since most Adapter-based methods have residual connections, we can multiply the adapter modules with $\alpha$ to control their strengths. A similar approach can be applied to efficient selective methods (e.g.LoRA) as they learn additive residuals to the original parameters. As shown in 1(b) (more results in Appendix C), WiSE improves both fine-tuning and distribution shift performance of PETL methods. Interestingly, even though full fine-tuning is generally less robust than PETL methods, applying WiSE allows it to achieve better performance in both target distribution and distribution shift data, which suggests a promising research direction for robust PETL.

### v4

@@ 2.2 Parameter-Efficient Fine-Tuning (PEFT)
Table 1: PEFT methods performances on VTAB-1K (19 tasks from 3 groups) by Top-1 Accuracy. Based on the results across PEFT, linear probing, and full FT, we identify two task groups (purple and orange), as discussed in section 6.
2.3 Related work and comparison
The community-wide enthusiasm for PEFT has led to multiple survey articles [101, 99, 28]. Meanwhile, several empirical and theoretical studies were presented, mostly based on NLP tasks, attempting to provide a holistic understanding. [29, 67] provided unified views to methodologically connect PEFT methods. [10, 17, 31] and [32, 98] empirically compared PEFT methods on NLP and vision tasks, respectively, while [22] offered a theoretical stability and generalization analysis. Accuracy-wise, [10, 17, 31] found that PEFT is robust to over-fitting and quite effective in NLP tasks under low-data regimes. This is, however, not the case for vision tasks: [32] showed that representative PEFT methods like LoRA and Adapter can’t consistently outperform either full FT or linear probing. In terms of why PEFT works, [22] framed PEFT as sparse fine-tuning and showed that it imposes a regularization by controlling stability; [17, 32] framed PEFT as (subspace) optimization; [17] further discussed the theoretical principle inspired by optimal control.
Our study strengthens and complements the above studies and offers new insights. First, we compared over ten PEFT methods and carefully tuned the hyper-parameters, aiming to accurately assess each method’s performance. This is particularly crucial for the vision community, where comprehensive references are still limited, and simpler methods like BitFit have often been deemed inferior, while other approaches have shown discrepancies compared to NLP studies. Second, we go beyond a competition perspective to investigate a complementary perspective of PEFT approaches. We show that different PEFT approaches offer effective base learners for model ensembles. Third, we go beyond downstream accuracy to investigate PEFT’s effectiveness in maintaining out-of-distribution robustness. Finally, we analyze both low-shot and many-shot settings, revealing distinct patterns among PEFT, full FT, and linear probing, extending the understanding of PEFT.
3 PEFT Methods in Low-Shots Regime
@@
Figure 2: Ranking frequency of 15 methods (14 PEFT + linear probing) for three groups in VTAB-1K. Element $(i,j)$ is the number of times method $i$ ranks $j^{th}$ in each group. Methods are ordered by mean ranks (in brackets). The parameters column shows the # of trainable parameters in millions. More details are in Appendix C.
@@
Pre-trained models are meant to ease downstream applications. One representative scenario is low-shot learning: supervised FT of the pre-trained model with a small number of samples per class. Indeed, low-shot learning has been widely used to evaluate PEFT performance.
@@
Dataset. VTAB-1K [104] consists of 19 classification tasks from three groups. Natural comprises natural images captured with standard cameras. Specialized contains images captured by specialist equipment for remote sensing and medical purposes. Structured evaluates scene structure comprehension, including object counting and depth estimation. Following [104], we split the 1000 training image 80/20 for hyperparameter tuning. The reported Top-1 Accuracy is obtained after training over the 1000 images and evaluating on the original test set.
@@
Methods. We consider linear probing, full FT, and 14 PEFT methods: 2 prompt-based [42], 5 adapter-based [36, 74, 11, 43, 63], 4 direct selective [103, 7, 97, 55], and 3 efficient selective [37, 44]. Please refer to subsection 2.2 and Appendix B for details.
@@
Setup. We employ the ViT-B/16 [19] pre-trained on ImageNet-21K [15] as the backbone. The prediction head is randomly initialized for each dataset. We systematically tune 1) learning rate, 2) weight decay, and 3) method-specifics like the PEFT parameter sizes, which are often left intact in previous studies. We set a cap for PEFT size $\leq 1.5\%$ of ViT-B/16. We also turn the drop path rate [38] on (0.1) or off (0). A detailed experiment setup is provided in Appendix A, and more experiment results, including DINOv2 [71] and larger backbones (ViT-L and ViT-H), are provided in Appendix C.
@@
Results. As shown in 1(a) and Table 1, PEFT methods generally outperform both linear probing and full FT across datasets. Additionally, with proper implementation and fair hyperparameter tuning, we surprisingly found that most PEFT methods perform similarly as the relative standard deviations (divided by the means) in all three groups are quite low. Simple methods (e.g., Bitfit) and PEFT methods originally proposed for NLP (e.g., LoRA and Adapter), which were previously reported as inferior due to unoptimized implementations and hyperparameter tuning, now demonstrate competitive performance with SOTA visual PEFT methods. To understand the relative advantages of different approaches, we provide the ranking frequency of PEFT methods across different groups in Figure 2, where the element $(i,j)$ in each ranking matrix represents the frequency that method $i$ ranks $j^{th}$ in each group. Methods are ordered by their mean ranks (in brackets), and the parameters column indicates the number of trainable parameters in millions. In natural group, simpler methods with fewer trainable parameters—such as DiffFit and Fact-TT—offer a cost-effective solution without compromising performance. Conversely, in specialized and structured groups, methods with more parameters generally yield better performance. We hypothesize that this performance discrepancy arises from the domain similarity between the pre-trained domain (ImageNet) and the downstream domains. The natural group, sharing a stronger affinity with ImageNet, allows simpler methods like BitFit to adjust the features effectively. In contrast, the specialized and structured groups necessitate more complex methods with more trainable parameters to bridge the domain gap.
@@
Recipes. In low-shot regimes, when the downstream data are similar to the pre-trained data, simple methods (e.g., DiffFit) offer solid accuracy with fewer parameters. Conversely, if there is a substantial domain gap, more complex methods with more parameters often achieve higher accuracy. Since low-shot training is especially prone to overfitting, we recommend activating a nonzero drop path rate—commonly set to zero by default—that stochastically drops a Transformer block per sample [38]. All methods can benefit from such a randomization-based regularization, as shown in Figure 11 in the Appendix.
4 Different PEFT Approaches Offer Complementary Information
The previous section demonstrates that all PEFT methods perform similarly across various domains. Given that different PEFT methods are trained on the same downstream data using the same backbone and achieve comparable accuracy, one might expect them to learn similar knowledge from the data, resulting in similar predictions. Contrary to this expectation, our findings below reveal that different PEFT methods acquire distinct and complementary knowledge from the same downstream data, even when built upon the same backbone, leading to diverse predictions.
Figure 3: (a) Prediction similarity analysis: element $(i,j)$ shows the percentage of samples that method $i$ and $j$ predict the same. Although different methods achieve similar accuracy, they have diverse predictions. (b) LoRA, Adapter, and SSF FT on CIFAR100 (VTAB-1K). Their wrong prediction overlaps for the 5K least confident data. Correct prediction overlaps for the 5K most confident data are shown in 1(b). More results are in Appendix C.
Figure 4: Ensemble (majority vote) shows consistent gain on most datasets thanks to the diverse predictions.
We start by analyzing their prediction similarity on the same dataset in VTAB-1K. It is expected that their predictions are similar for datasets with very high accuracy, such as Flowers102 (avg 99.1%) and Caltech101 (avg 91.4%). Beyond them, we find that most PEFT methods show diverse predictions in other datasets in VTAB-1K. 3(a) shows the prediction similarities between 14 PEFT methods in DTD, Retinopathy, and DMLab, which belong to natural, specialized, and structured groups, respectively. In DTD and Retinopathy, most methods differ in about 20% of their predictions, while in DMLab, this difference increases to approximately 35%, even though they achieve similar accuracies. This prediction diversity may be attributed to the different inductive biases [70] of PEFT methods — they explicitly select specific parameters to update or insert different modules at various locations within the model. More analyses and details are offered in Appendix C.
Such diverse predictions across methods open up the possibility of leveraging their heterogeneity for further improvement. The most straightforward approach is ensemble [26], e.g., majority vote over methods. Figure 4 demonstrates the ensemble performance gain over all the PEFT methods in each dataset, where we use the worst PEFT method as the baseline. Thanks to the diverse predictions across methods, the ensemble can provide consistent gain.
Also, we analyze if PEFT methods make similar correct predictions for high-confidence samples and similar mistakes for low-confidence samples. 1(b) and 3(b) show the correct prediction overlap for the 5K most confident samples (per method) and the wrong prediction overlap for the 5K least confident samples (per method). For demonstration purposes, we select one method from each PEFT category (LoRA, Adapter, SSF) and they are FT on CIFAR-100 in VTAB-1K. Methods within the same category also show diverse predictions (Appendix C). Since they make different predictions in both high and low-confidence regimes, this paves the way for new possibilities of using different PEFT methods to generate diverse pseudo-labels for semi-supervised learning [100, 23, 24], domain adaptation [21, 20, 86], and continual learning [66, 65, 82, 60, 64]. For example, in semi-supervised learning, we can FT different PEFT methods on the labeled data to generate diverse and accurate pseudo-labels by selecting highly confident predictions from each PEFT method.
5 PEFT Methods in Many-Shot Regime
Recent works in NLP [10] have indicated that PEFT methods may not perform as competitively as full FT when data is abundant. We thus aim to investigate PEFT’s performance in many-shot regimes by addressing the following questions: (1) Should we use PEFT or full FT when data is sufficient? (2) How should we adjust the number of trainable parameters for PEFT methods in many-shot regimes?
Dataset. We select one representative dataset from each group in VTAB: (1) CIFAR-100 [49], a natural image dataset comprising 50K training images across 100 classes; (2) RESISC [12], a remote sensing dataset for scene classification with 25.2K training samples across 45 classes; and (3) Clevr-Distance [104, 45], a synthetic image dataset for predicting the depth of the closest object with 6 depth classes and 70K samples. The reported results are obtained by training on the full training set and evaluating on the original test set.
Setup. The model setup mostly follows the VTAB-1K experiment. More details about setup and hyperparameter search are provided in Appendix A.
Results. In many-shot regimes, with sufficient downstream data, full FT may catch up and eventually outperform PEFT methods. However, from Figure 5, we found that even in many-shot regimes, PEFT can achieve comparable results with full FT, even just using $2\%$ of fine-tuning parameters. The performance gain, however, quickly diminishes and plateaus after 5% of tunable parameters. By comparing the results on the domain-close CIFAR-100 and domain-different RESISC and Clevr, we have some further observations. Downstream tasks with larger domain gaps often require updating more parameters to achieve high accuracy. With sufficient downstream data, full FT is less prone to over-fitting and, indeed, attains a high accuracy. But interestingly, PEFT methods, with only 2$\sim$5of tunable parameters, achieve similar accuracy, suggesting that its design principle does offer sufficient effective capacity for the model to learn [105]. Downstream tasks with smaller domain gaps suggest that the pre-trained model had learned sufficient knowledge about them; fully FT it thus risks washing such knowledge away. In fact, we found that PEFT notably outperforms full FT on CIFAR-100, suggesting it as a more robust transfer learning algorithm for downstream tasks.
Recipes. In many-shot regimes, PEFT methods with sufficient parameters (2$\sim$5%) appear more favorable than full FT and linear probing. On the one hand, they achieve comparable and even better accuracy than full FT. On the other hand, the tunable parameters remain manageable. The parameter efficiency of PEFT also often implies less training GPU memory usage and training time, making PEFT methods a favorable alternative in many-shot regimes. For a downstream domain that is close to the pre-training domain, PEFT shows much pronounced transferability. For a downstream domain that is quite different, the limited tunable parameters (controversially, 2$\sim$5already amount to a few million) already allow the model to learn sufficiently.
@@ 5 PEFT Methods in Many-Shot Regime
Figure 5: PEFT accuracy in many-shot regimes, with different parameter sizes (X-axis) on three datasets from different domains. Even 2%-5% trainable parameters allow the models to have sufficient capacity to learn from full data. Details are in Appendix C).
6 Why Do PEFT Methods Work?22footnotemark: 2
footnotetext: Our intention is not to offer a definitive explanation about why PEFT works. As discussed in subsection 2.3, there is no broadly accepted theory for PEFT’s success. We hope our empirical findings can support the ongoing efforts to uncover the fundamental principles behind PEFT.
@@ 6 Why Do PEFT Methods Work?22footnotemark: 2
Putting together section 3 and section 5, we identify two distinct patterns regarding the performance among linear probing, full FT, and PEFT. Within 19 VTAB-1K tasks, we see: (1) Full FT outperforms linear probing. As linear probing reflects the pre-trained feature quality for downstream tasks, case (1) suggests the necessity to update the backbone to close the gap between pre-trained and downstream domains. (2) Linear probing surpasses full FT, suggesting the pre-trained features are good enough (at least in a low-shot scenario). Recklessly updating them may risk over-fitting. Figure 6 (a-b) summarizes the low-shot accuracy comparison based on the categorization above; each line corresponds to one task. Linear probing, PEFT, and full FT are located in order, from left to right, to reflect their tunable parameter sizes. PEFT’s superiority in both cases showcases its capacity to learn and its regularization role to prevent over-fitting.
@@
We also draw the many-shot accuracy in Figure 6 (c-d) based on the same categorization: RESISC and Clevr in case (1), and CIFAR-100 in case (2). In the many-shot setting, full FT consistently outperforms linear probing, which seems to suggest no more risk of over-fitting. However, on CIFAR-100 (Figure 6 (d)), we again see a noticeable gap between PEFT and full FT, just like in Figure 6 (b). Such a concave shape reminds us of the long-standing under-fitting-over-fitting curve, suggesting that even with sufficient downstream data, full FT still risks over-fitting.
@@
Considering PEFT’s comparable performance to full FT on RESISC and Clevr with large domain gaps, we conclude that PEFT succeeds as a high-capacity learner equipped with an effective regularizer. The two roles trade-off well such that PEFT can excel in both low- and high-similarity domains under both low-shot and many-shot settings.
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
@@
Figure 6: (a): VTAB-1K tasks in case (1), PEFT $>$ full $>$ linear. (b) VTAB-1K tasks in case (2), PEFT $>$ linear $>$ full. (c) RESISC & Clevr in case (1) with enough data, PEFT $\approx$ full $>$ linear. (d) CIFAR in case (2) with enough data, PEFT $>$ full $>$ linear. Within each figure, left for linear, middle for PEFT, and right for full. More details are in Appendix C.
@@
7 How Robust are PEFT Methods to Distribution Shifts?
Large pre-trained models such as CLIP [75] have demonstrated unprecedented zero-shot accuracy across diverse data distributions. However, recent studies [96, 75] have shown that FT on downstream data, while significantly boosting performance on the target distribution, often compromises the model’s robustness to distribution shifts. Given that PEFT only updates a limited number of parameters in the model, we investigate whether PEFT can offer a more robust alternative to full FT.
Dataset. We use 100-shot ImageNet-1K as our target distribution, with each class containing 100 images. Following [96], we consider 4 natural distribution shifts from ImageNet: ImageNet-V2 [78], a new ImageNet test set collected with the original labeling protocol; ImageNet-R [34], renditions for 200 ImageNet classes; ImageNet-S [25], sketch images for 1K ImageNet classes; ImageNet-A [35], a test set of natural images misclassified by a ImageNet pre-trained ResNet-50 [30] for 200 ImageNet classes.
Setup. We focus on the CLIP ViT-B/16 model, which comprises a visual encoder and a text encoder, pre-trained via contrastive learning on image-text pairs. Following [96], we add an FC layer as the head initialized using the class label text embedded by the text encoder. Subsequently, we discard the text encoder and apply PEFT methods to the visual encoder, FT only the PEFT modules and the head. More details about the CLIP model and experiment setup can be found in Appendix A.
Results. As shown in Table 2, while some PEFT methods may not surpass full fine-tuning on the target distribution, they consistently demonstrate more robust performance on distribution-shifted data. This robustness can be attributed to PEFT updating only a small fraction of the parameters, thereby preserving the robust features of the foundational models. Given the similar target distribution performance, should we blindly use PEFT methods for more robustness?
Weight-space ensembles (WiSE) for PEFT. WiSE [96], which linearly interpolates the weights of a fully FT model with those of the original backbone, is a popular approach for boosting robustness. We explore whether WiSE can also enhance the robustness of PEFT. To apply WiSE to PEFT, we first linearly interpolate the prediction head with a mixing coefficient $\alpha$. For direct selective tuning methods (e.g. BitFit), this involves merging the PEFT-tuned parameters and the original model. Since most Adapter-based methods include residual connections (subsubsection B.2.2), we can adjust their impact by scaling the adapter modules with $\alpha$. A similar approach applies to efficient selective methods (e.g. LoRA) as they learn additive residuals to the original parameters. To the best of our knowledge, we are the first to study WiSE for PEFT. As shown in 1(c) (more results in Appendix C), WiSE consistently improves both target and distribution shift performances of PEFT methods. For Adapter-based methods, WiSE can be considered feature ensembles, where $\alpha$ controls how strong the domain-specific features from the adapter module blend with the domain-agnostic ones from original backbones. For selective tuning, WiSE functions similarly to its application in full FT—exploiting the fact that the fine-tuned parameters remain near the original loss basin, allowing an ensemble that captures the best of both [2, 40].
Interestingly, even though full FT is generally less robust than PEFT methods, WiSE elevates full FT’s performance above that of PEFT on both target distribution and distribution shift data, which suggests promising research directions to investigate the underlying mechanisms and how to further improve the robustness of PEFT methods.